EQUILIBRIUM STRUCTURE OF THE SIMPLE SKEW CHAIN MOLECULE HSOH

\underline{OLIVER BAUM}, STEPHAN SCHLEMMER, THOMAS F. GIESEN, I. Physikalisches Institut, Universität zu Köln, 50937 Köln, Germany; JÜRGEN GAUSS, Institut für Physikalische Chemie, Universität Mainz, 55099 Mainz, Germany.

In this paper we present an empirical equilibrium structure of HSOH, which is either known as oxadisulfane or hydrogen thioperoxide. The molecule is non planar with a simple skew chain structure analogues to the kin molecules hydrogenperoxide, HOOH, and hydrogen-disulfane, HSSH. Gas-phase studies on oxadisulfane were long time hampered due to the problems of synthesizing the molecule. As recently as in 2001 Behnke^{a} found a way to produce the molecule by pyrolysis of di-tert-butyl sulfoxide. The first high resolution gas-phase measurements on HSOH were presented in 2003 by Winnewisser et al.^{b}. Beside the measurements on HSOH they were able to assign the spectrum of H^{34}SOH in natural abundance. In 2003 Behnke et al.^{c} also recorded the spectra perdeutero oxadisulfane, DSOD, which was produced in a discharge of D_{2}S together with D_{2}O. Recently Baum et al.^{d} reported the first high resolution measurements on single deuterated oxadisulfane, HSOD. The molecule was synthesized in a rf-discharge of pure sulphur and HDO. Based on these highly precise data from our measurements on HSOH, H^{34}SOH, DSOD^{e}, and HSOD we derived an empirical equilibrium structure of oxadisulfane. The equilibrium rotational constants A_{e}, B_{e}, and C_{e} were obtained from A_{0}, B_{0}, and C_{0} which were corrected by vibration-rotation interactions constants α_{r} from high-level ab initio calculations. The derived bond lengths and angles agree very well with those from high level quantum chemical calculations.

aM. Behnke, Ph.D. thesis, Universität zu Köln (2001)

bG. Winnewisser, F. Lewen, S. Thorwirth, M. Behnke, J. Hahn, J. Gauss, E. Herbst, J. Chem. Phys. 9 (2003) 5501–5510

cM. Behnke, J. Suhr, S. Thorwirth, F. Lewen, H. Lichau, J. Hahn, J. Gauss, K. M. T. Yamada, G. Winnewisser, J. Mol. Spectrosc.

221 (2003) 121–126

dO. Baum, S. Esser, N. Gierse, S. Brünken, F. Lewen, J. Hahn, J. Gauss, S. Schlemmer, T. F. Giesen, J. Mol. Struct. (2006)

(accepted)

eS. Brünken, M. Behnke, S. Thorwirth, K. M. T. Yamada, T. F. Giesen, F. Lewen, J. Hahn, G. Winnewisser, J. Mol. Struct.

742 (2005) 237–242